The length of a rectangle is 3 in. more than its width. If the length were decreased by 2 in. and the width were increased by 1 in., the perimeter of the resulting rectangle would be 24 in. Find the dimensions of the original rectangle.

Answer by lwsshak3(11628)

The length of a rectangle is 3 in. more than its width. If the length were decreased by 2 in. and the width were increased by 1 in., the perimeter of the resulting rectangle would be 24 in. Find the dimensions of the original rectangle.
***
let x=width of original rectangle
x+3=length of original rectangle
..
x+3-2=x+1 ( length decreased by 2 in)
x+1 (width increased by 1 in)
perimeter=twice the width+twice the length
2(x+1)+2(x+1)=24
2x+2+2x+2=24
4x=20
x=5
x+3=8
width of original rectangle=5 in
length of original rectangle=8 in